# NPGA: Neural Parametric Gaussian Avatars

### v1

@@ Abstract.
The creation of high-fidelity, digital versions of human heads is an important stepping stone in the process of further integrating virtual components into our everyday lives. Constructing such avatars is a challenging research problem, due to a high demand for photo-realism and real-time rendering performance. In this work, we propose Neural Parametric Gaussian Avatars (NPGA), a data-driven approach to create high-fidelity, controllable avatars from multi-view video recordings. We build our method around 3D Gaussian splatting for its highly efficient rendering and to inherit the topological flexibility of point clouds. In contrast to previous work, we condition our avatars’ dynamics on the rich expression space of neural parametric head models (NPHM), instead of mesh-based 3DMMs. To this end, we distill the backward deformation field of our underlying NPHM into forward deformations which are compatible with rasterization-based rendering. All remaining fine-scale, expression-dependent details are learned from the multi-view videos. To increase the representational capacity of our avatars, we augment the canonical Gaussian point cloud using per-primitive latent features which govern its dynamic behavior. To regularize this increased dynamic expressivity, we propose Laplacian terms on the latent features and predicted dynamics. We evaluate our method on the public NeRSemble dataset, demonstrating that NPGA significantly outperforms the previous state-of-the-art avatars on the self-reenactment task by $\approx 2.6$ PSNR. Furthermore, we demonstrate accurate animation capabilities from real-world monocular videos.
footnotetext: Project Website: https://simongiebenhain.github.io/NPGA/
Figure 1. NPGA: We utilize the rich expression space of Neural Parametric Head Models to create high-fidelity avatars with fine-grained expression control. Our avatars consist of a dynamics module and a canonical Gaussian point cloud, which is augmented with per-primitive features that encode valuable semantic information, as indicated on the left. On the right, we demonstrate a highly detailed cross-reenactment using the inset image as a driving expression.
@@ 1. Introduction
Creating photo-realistic 3D avatars is one of the core challenges in computer graphics and includes a wide range of applications such as movies, games, AR/VR teleconferencing, and the metaverse. In particular, there is a strong motivation to reconstruct digital avatars from real-world captures, such as multi-view recordings, to obtain a digital copy of a specific real person. The resulting digital avatars can then be animated and rendered from arbitrary viewpoints while expecting high visual fidelity; e.g., with respect to photo-realistic colors and details, preservation of identity, and the adoption of person-specific mannerisms. At the same time, many avatar applications demand real-time rendering capabilities without compromising visual quality.
@@
To this end, we propose NPGA, a new avatar representation that leverages a learned deformation representation while ensuring that the predicted facial dynamics stay close to the prior of an underlying neural parametric head model (NPHM) (Giebenhain et al., 2023, 2024). NPHM provides our avatars with more fine-grained expression control compared to classical, public 3DMMs (Paysan et al., 2009; Li et al., 2017), which were previously used for avatar creation. NPGA consists of a canonical Gaussian point cloud, that can be forward-deformed using an expression code and rendered using 3DGS, similar to previous work (Zielonka et al., 2023; Qian et al., 2023; Xu et al., 2024). As a rasterization-based approach, 3DGS cannot be efficiently combined with the backward deformation field of MonoNPHM. Therefore, we propose a distillation strategy to invert the deformation direction of MonoNPHM’s expression prior using a cycle consistency loss. The resulting forward deformation field becomes compatible with the rasterization-based rendering of 3DGS. To increase the overall dynamic expressivity of our avatars, we further propose to augment our canonical Gaussian with per-Gaussian latent features, which enable our deformation module to operate in a higher dimensional space, that can describe facial dynamics more effectively. We show that this added expressivity results in higher-fidelity image synthesis, but is required to be appropriately regularized to achieve artifact-free renderings. To this end, we formulate Laplacian smoothness terms on the latent features and predicted dynamics, based on the k-nearest neighbor graph in canonical space. Furthermore, we modify the adaptive density control strategy of 3DGS for more detailed avatar reconstructions.
@@
  • We propose a distillation strategy to utilize NPHM’s rich expression prior for more expressive and controllable avatars.
  • We introduce per-Gaussian latent features that lift the input domain of our deformation module to a higher-dimensional space, yielding increased dynamic capacity.
  • Our avatars outperform the previous state-of-the-art by 2.6 PSNR and 0.021 SSIM on the self-reenactment task. Furthermore, we demonstrate accurate avatar reconstruction from monocular RGB sequences.
@@ 2.1. Dynamic Scene Representations
Traditional morphable models (Blanz and Vetter, 1999; Paysan et al., 2009) learn a representation of body geometry via PCA. They are one of the primary tools to drive human animation and are heavily used in industry applications, making them a core building block for work on virtual avatars. While some models are dedicated to specific regions like the face (Paysan et al., 2009) and head (Li et al., 2017), some variants include the neck (Zhang et al., 2023) or even the entire body (Pavlakos et al., 2019). More recently, neural equivalents of 3DMMs, such as i3DMM (Yenamandra et al., 2021), ImFace (Zheng et al., 2022a) or NPHM (Giebenhain et al., 2023), have improved upon the expression fidelity compared to classical PCA-based models. They encode a canonical representation of the geometry via a signed distance functions (SDF), which can be mapped to arbitrary expressions via a deformation field. NPGA utilizes NPHM as it offers several beneficial characteristics: It models the face densely including eyes, hair, and teeth, it captures local details well and it disentangles shape from expressions.
@@ 2.3. Human Head Reconstruction and Animation
Existing approaches for animating avatars mainly differ in two fundamental aspects: first, the utilized 3D representation in combination with its rendering mechanism, and second, how the expression codes are transferred into scene dynamics. Existing work has explored, among others, meshes in combination with deferred neural rendering (Grassal et al., 2022; Kim et al., 2018), neural radiance fields (Gafni et al., 2021; Athar et al., 2022; Zielonka et al., 2022), and 3D CNNs in combination with volume rendering (Lombardi et al., 2019, 2021; Cao et al., 2022). Recently, there has been a lot of interest in point-based representations and rendering (Zheng et al., 2023; Qian et al., 2023; Xu et al., 2024), especially since Kerbl et al. (2023a) proposed 3D Gaussian Splatting, which we also adopt in our work, due to its efficient rendering and topological flexibility. While some approaches choose to explain the face’s movement explicitly through the underlying mesh of a 3DMM, e.g., (Athar et al., 2022; Zielonka et al., 2022; Qian et al., 2023), others choose the opposing extreme of a more data-driven approach by freely learning the face movement using a neural component, which is directly conditioned on the expression codes (Gafni et al., 2021; Xu et al., 2024; Lombardi et al., 2021). NPGA adopts the latter idea but replaces the 3DMM with a neural parametric model.
@@ 4. Method
Fig. 2 shows an overview of our proposed representation and methodology to build our Neural Parametric Gaussian Avatars (NPGA). As described in Section 4.1, our avatars are composed of two key components: a canonical Gaussian point cloud $\mathcal{A}_{c}$ and a dynamics module $\mathcal{D}$ which deforms the Gaussians when provided with an expression code, similar to recent work  (Qian et al., 2023; Xu et al., 2024). In Section 4.2 we describe our distillation strategy that allows NPGA to leverage the rich latent expression space and detailed motion prior of MonoNPHM (Giebenhain et al., 2024). Given multi-view video recordings alongside tracked MonoNPHM expression codes, we jointly optimize for our canonical Gaussians and dynamics module, as described in Section 4.3.
@@ 4.1.1. Canonical Representation
with per-Gaussian features $\mathbf{f}\in\mathbb{R}^{N\times 8}$. While these features are static themselves, they provide crucial semantic information to describe the dynamic behavior of the respective primitives. In some sense, our per-Gaussian features serve a similar purpose as positional encodings (Mildenhall et al., 2021; Müller et al., 2022), which are uncorrelated with spatial coordinates, have infinite spatial resolution and do not require additional data structures.
@@ 4.1.3. Screen-Space Refinement
where the $\hat{I}_{\text{rgb}}$ denotes rendered RGB colors. $I_{h}$ is a latent image used for the CNN refinement module, which is obtained by rendering $\mathbf{h}+\delta_{\mathbf{h}}^{\mathcal{G}}$, where $\mathbf{h}$ are latent CNN features which we additionally include in $\mathcal{A}_{c}$. Note, that compared to (Xu et al., 2024) we do not perform super-resolution, but motivate the use of the CNN through increased performance in our ablation study.
@@ 4.2. Cycle-Consistency Distillation
One of our core ideas is to leverage MonoNPHM’s motion prior and expression space. However, since it utilizes a backward deformation field $\mathcal{B}$, which warps points into canonical space, we cannot directly incorporate this deformation prior in our pipeline. Instead, we need forward deformations, which warp points into posed space, such that they can be directly rasterized. While it is possible to numerically approximate the inverse of $\mathcal{B}$ using iterative root-finding (Chen et al., 2021, 2023), we search for a more computationally efficient method. Instead, we propose to distill a forward deformation network $\mathcal{F}$ as the inverse of $\mathcal{B}$ using a cycle consistency loss
@@
represented as low-resolution 64x64 triplanes (Peng et al., 2020; Chan et al., 2022), which can be evaluated at arbitrary points $x_{c}$ in canonical space.
We train $\mathcal{F}$ once, using the tracked sequences of 20 people in the NeRSemble dataset (Kirschstein et al., 2023b), and use the same $\mathcal{F}$ for all our avatars. During training, each person has their own TriPlane, which we regularize using a total variation loss. Furthermore, we regularize the norm of predicted offsets $\|\mathcal{F}(x_{c},\mathbf{f}(x_{c})\|_{2}$ to be small.
@@ 4.3. Avatar Optimization Strategy
After obtaining a forward deformation field $\mathcal{F}$ using our cycle-consistency distillation strategy, we aim to jointly optimize for the canonical parameters $\mathcal{A}_{c}$ and MLP $\mathcal{G}$ to minimize a photometric energy term. To initialize the canonical Gaussians centers $\mu$ we sample $30.000$ points uniformly on the iso-surface of the tracked MonoNPHM model. The per Gaussian features are initialized by querying $\textsc{TriPlane}(\mu)$ at the sampled Gaussian centers. All remaining attributes are initialized using the default 3DGS procedure. In practice, we observed that keeping $\mathcal{F}$ frozen results in sub-optimal performance, which is likely caused through topological issues during distillation in the mouth region. Hence, we decide to further optimize $\mathcal{F}$ alongside $\mathcal{G}$ and $\mathcal{A}_{c}$, however, using a significantly smaller learning rate and a warm-up schedule that encourages the preservation of the distilled prior.
@@ 5.1.1. Baselines
5.1.2. Metrics
Driving Expression MVP GaussianAvatars GHA Ours
5.2. Implementation Details
Hyper-Parameters
Runtime
While we do not focus on efficient training, animation, or rendering of avatars, we acknowledge the importance of fast animation and rendering. In our unoptimized implementation, we can render images at 31 frames per second (FPS) for 550x802 and 18 FPS at 1100x1604 on an NVIDIA RTX3080 graphics card, which includes deformation, rendering, and CNN. When omitting the CNN the speed increases to 43 and 38 FPS, respectively. As a comparison, GHA runs at 22 FPS at 1024x1024 on the same machine. We train all our avatars, and baselines, until convergence, which roughly takes 7 hours for GA (on an RTX2080), 30 hours (on an RTX3090) for GHA and our method, and 60 hours (on an RTX2080) for MVP.
Data Preparation
@@ 5.2. Implementation Details
Ours 37,68 0,973 0,032 30.42 0.935 0.057

### v2

@@ Abstract.
The creation of high-fidelity, digital versions of human heads is an important stepping stone in the process of further integrating virtual components into our everyday lives. Constructing such avatars is a challenging research problem, due to a high demand for photo-realism and real-time rendering performance. In this work, we propose Neural Parametric Gaussian Avatars (NPGA), a data-driven approach to create high-fidelity, controllable avatars from multi-view video recordings. We build our method around 3D Gaussian splatting for its highly efficient rendering and to inherit the topological flexibility of point clouds. In contrast to previous work, we condition our avatars’ dynamics on the rich expression space of neural parametric head models (NPHM), instead of mesh-based 3DMMs. To this end, we distill the backward deformation field of our underlying NPHM into forward deformations which are compatible with rasterization-based rendering. All remaining fine-scale, expression-dependent details are learned from the multi-view videos. For increased representational capacity of our avatars, we propose per-Gaussian latent features that condition each primitives dynamic behavior. To regularize this increased dynamic expressivity, we propose Laplacian terms on the latent features and predicted dynamics. We evaluate our method on the public NeRSemble dataset, demonstrating that NPGA significantly outperforms the previous state-of-the-art avatars on the self-reenactment task by $\approx 2.6$ PSNR. Furthermore, we demonstrate accurate animation capabilities from real-world monocular videos.
Virtual avatars, 3D Gaussian splatting, Data-driven animation, 3d morphable models
journal: TOGjournalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687689isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Trackingccs: Computing methodologies Motion capture
@@ 1. Introduction
Creating footnotetext: Project Website: https://simongiebenhain.github.io/NPGA/ photo-realistic 3D avatars is one of the core challenges in computer graphics and includes a wide range of applications such as movies, games, AR/VR teleconferencing, and the metaverse. In particular, there is a strong motivation to reconstruct digital avatars from real-world captures, such as multi-view recordings, to obtain a digital copy of a specific real person. The resulting digital avatars can then be animated and rendered from arbitrary viewpoints while expecting high visual fidelity; e.g., with respect to photo-realistic colors and details, preservation of identity, and the adoption of person-specific mannerisms. At the same time, many avatar applications demand real-time rendering capabilities without compromising visual quality.
@@
To this end, we propose NPGA, a new avatar representation that leverages a learned deformation representation while ensuring that the predicted facial dynamics stay close to the prior of an underlying neural parametric head model (NPHM) (Giebenhain et al., 2023, 2024). NPHM provides our avatars with more fine-grained expression control compared to classical, public 3DMMs (Paysan et al., 2009; Li et al., 2017), which were previously used for avatar creation. In another line of work Cao et al. (2022) explore a learned neural prior over faces, that helps to map baked uv-space information from a monocular video into a volumetric avatar representation.
NPGA consists of a canonical Gaussian point cloud, that can be forward-deformed using an expression code and rendered using 3DGS, similar to previous work (Zielonka et al., 2023; Qian et al., 2023; Xu et al., 2024). As a rasterization-based approach, 3DGS cannot be efficiently combined with the backward deformation field of MonoNPHM. Therefore, we propose a distillation strategy to invert the deformation direction of MonoNPHM’s expression prior using a cycle consistency loss, similar to SCANimate (Saito et al., 2021). The resulting forward deformation field becomes compatible with the rasterization-based rendering of 3DGS. To increase the overall dynamic expressivity of our avatars, we further propose to augment our canonical Gaussian with per-Gaussian latent features similar to GaussianHeadAvatar (Xu et al., 2024) and HeadGAS (Dhamo et al., 2023). Compared to these works, we allow these features to influence the movement of Gaussians instead of only influencing dynamic changes in appearance. This enables our deformation module to operate in a higher dimensional space, that can describe facial dynamics more effectively. We show that this added expressivity results in higher-fidelity image synthesis, but is required to be appropriately regularized to achieve artifact-free renderings. To this end, we formulate Laplacian smoothness terms on the latent features and predicted dynamics, based on the k-nearest neighbor graph in canonical space. Furthermore, we modify the adaptive density control (ADC) strategy of 3DGS for more detailed avatar reconstructions.
@@
  • To leverage MonoNPHM’s rich expression prior for fine-grained animation control and effective optimization, we utilize a cycle-consistency-based distillation strategy.
  • We condition our deformation network on per-Gaussian latent features for increased dynamic capacity, use Laplacian regularization, and adjust the ADC for a dynamic setting.
  • We outperform the previous state-of-the-art by 2.6 PSNR and 0.021 SSIM. Furthermore, we demonstrate accurate avatar re-animation from monocular RGB sequences.
@@ 2.1. Dynamic Scene Representations
Traditional morphable models (Blanz and Vetter, 1999; Paysan et al., 2009) learn a representation of body geometry via PCA. They are one of the primary tools to drive human animation and are heavily used in industry applications, making them a core building block for work on virtual avatars. While some models are dedicated to specific regions like the face (Paysan et al., 2009) and head (Li et al., 2017), some variants include the neck (Zhang et al., 2023) or even the entire body (Pavlakos et al., 2019). More recently, neural equivalents of 3DMMs, such as i3DMM (Yenamandra et al., 2021), ImFace (Zheng et al., 2022a) or NPHM (Giebenhain et al., 2023, 2024), have improved upon the expression fidelity compared to classical PCA-based models. They encode a canonical representation of the geometry via a signed distance functions (SDF), which can be mapped to arbitrary expressions via a deformation field. NPGA utilizes NPHM as it offers several beneficial characteristics: It models the face densely including eyes, hair, and teeth, it captures local details well and it disentangles shape from expressions.
Instead of replacing mesh-based 3DMMs altogether, several approaches have been proposed to model details on top of an underlying PCA-driven mesh (Cao et al., 2022; Yang et al., 2020; Wang et al., 2022). Specifically, Cao et al. (2022) leverage high-frequency information from tracked UV texture maps to create and animate avatars, using a high-detailed, learned expression space.
@@ 2.3. Human Head Reconstruction and Animation
Existing approaches for animating avatars mainly differ in two fundamental aspects: first, the utilized 3D representation in combination with its rendering mechanism, and second, how the expression codes are transferred into scene dynamics. Existing work has explored, among others, meshes in combination with deferred neural rendering (Grassal et al., 2022; Kim et al., 2018), neural radiance fields (Gafni et al., 2021; Zhao et al., 2023; Zielonka et al., 2022), and CNNs in combination with primitive-based volume rendering (Lombardi et al., 2019, 2021; Cao et al., 2022). Recently, there has been a lot of interest in point-based representations and rendering (Zheng et al., 2023; Qian et al., 2023; Xu et al., 2024; Dhamo et al., 2023), especially since Kerbl et al. (2023a) proposed 3D Gaussian Splatting.While some approaches choose to explain the face’s movement explicitly through the underlying mesh of a 3DMM, e.g., (Athar et al., 2022; Zielonka et al., 2022; Qian et al., 2023), others choose the opposing extreme of a more data-driven approach by freely learning the face movement using a neural component, which is directly conditioned on the expression codes (Gafni et al., 2021; Xu et al., 2024; Lombardi et al., 2021). For NPGA we adopt the latter idea but replace the 3DMM with a neural parametric model, anticipating that the improved tracking quality translates to higher fidelity avatar reconstructions and animations.
@@ 4. Method
Fig. 2 shows an overview of our proposed representation and methodology to build our Neural Parametric Gaussian Avatars (NPGA). As described in Section 4.1, our avatars are composed of two key components: a canonical Gaussian point cloud $\mathcal{A}_{c}$ and a dynamics module $\mathcal{D}$ which deforms the Gaussians when provided with an expression code, similar to recent work  (Qian et al., 2023; Xu et al., 2024; Dhamo et al., 2023). In Section 4.2 we describe our distillation strategy that allows NPGA to leverage the rich latent expression space and detailed motion prior of MonoNPHM (Giebenhain et al., 2024). Given multi-view video recordings alongside tracked MonoNPHM expression codes, we jointly optimize for our canonical Gaussians and dynamics module, as described in Section 4.3.
@@ 4.1.1. Canonical Representation
with per-Gaussian features $\mathbf{f}\in\mathbb{R}^{N\times 8}$. While these features are static themselves, they provide crucial semantic information to describe the dynamic behavior of the respective primitives. In some sense, our per-Gaussian features serve a similar purpose as positional encodings (Mildenhall et al., 2021; Müller et al., 2022), which are uncorrelated with spatial coordinates, have infinite spatial resolution and do not require additional data structures. The idea of per-Gaussian features has been previously proposed in (Xu et al., 2024; Dhamo et al., 2023). Crucially, however, both works do not utilize these features to better predict the movement of Gaussians, which we ablate in Section 5.5. (Dhamo et al., 2023) goes beyond static per-primitive features by dynamically blending them with expression codes. In the formulation of (Xu et al., 2024) and ours, this happens implicitly inside the MLP introduced in the next paragraph.
@@ 4.1.3. Screen-Space Refinement
where the $\hat{I}_{\text{rgb}}$ denotes rendered RGB colors. $I_{h}$ is a latent image used for the CNN refinement module, which is obtained by rendering $\mathbf{h}+\delta_{\mathbf{h}}^{\mathcal{G}}$, where $\mathbf{h}$ are latent CNN features which we additionally include in $\mathcal{A}_{c}$. We took inspiration from Xu et al. (2024), to include a screen-space CNN, but note that we decided against performing super-resolution.
@@ 4.2. Cycle-Consistency Distillation
One of our core ideas is to leverage MonoNPHM’s motion prior and expression space. However, since it utilizes a backward deformation field $\mathcal{B}$, which warps points into canonical space, we cannot directly incorporate this deformation prior in our pipeline. Instead, we need forward deformations, which warp points into posed space, such that they can be directly rasterized. While it is possible to numerically approximate the inverse of $\mathcal{B}$ using iterative root-finding (Chen et al., 2021, 2023), we search for a more computationally efficient method. To this end we resort to a cycle consistency loss, similar to (Saito et al., 2021) and follow-ups (Dhamo et al., 2023; Yang et al., 2022) which optimize for volumetric skinning weights used for animation via linear blend skinning.
We propose to directly distill a forward deformation network $\mathcal{F}$ as the inverse of $\mathcal{B}$ (see Fig. 2, using a cycle consistency loss
@@
represented as low-resolution 64x64 triplanes (Peng et al., 2020; Chan et al., 2022), which can be evaluated at arbitrary points $x_{c}$ in canonical space. For convenience, we train $\mathcal{F}$ once, using all six identities from the NeRSemble dataset (Kirschstein et al., 2023b) that we perform experiments with.
During training, each person has their own TriPlane, which we regularize using a total variation loss. Furthermore, we regularize the norm of predicted offsets $\|\mathcal{F}(x_{c},\mathbf{f}(x_{c})\|_{2}$ to be small. Note that distilling $\mathcal{F}$ separately results in insignificant performance changes.
@@ 4.3. Avatar Optimization Strategy
After obtaining a forward deformation field $\mathcal{F}$ using our cycle-consistency distillation strategy, we aim to jointly optimize for the canonical parameters $\mathcal{A}_{c}$ and MLP $\mathcal{G}$ to minimize a photometric energy term. To initialize the canonical Gaussians centers $\mu$ we sample $50.000$ points uniformly on the iso-surface of the tracked MonoNPHM model. The per Gaussian features are initialized by querying $\textsc{TriPlane}(\mu)$ at the sampled Gaussian centers. All remaining attributes are initialized using the default 3DGS procedure. In practice, we observed that keeping $\mathcal{F}$ frozen results in sub-optimal performance, which is likely caused through topological issues during distillation in the mouth region. Hence, we decide to further optimize $\mathcal{F}$ alongside $\mathcal{G}$ and $\mathcal{A}_{c}$, however, using a significantly smaller learning rate and a warm-up schedule that encourages the preservation of the distilled prior.
@@ 4.3.2. Adaptive Density Control (ADC)
4.4. Differences to GaussianHeadAvatar (GHA)
In general our avatar formulation shares large parts with that of GHA (Xu et al., 2024). A major difference is our use of MonoNPHM as underlying 3DMM, which we enabled through our distillation strategy. Furthermore, we utilize the per-Gaussian features to also influence the prediction of $\delta_{\mu}$, we add a Laplacian regularization, do not rely on screen-space super resolution or perceptual losses, and modify the ADC strategy instead of resorting to a fixed number of Gaussians as Xu et al. (2024).
@@ 5.1.1. Baselines
[…]
Driving Expression MVP GaussianAvatars GHA Ours
5.1.2. Metrics
5.2. Implementation Details
Hyper-Parameters
@@ 5.2. Implementation Details
Runtime
While we do not focus on efficient training, animation, or rendering of avatars, we acknowledge the importance of fast animation and rendering. In our unoptimized implementation, we can render images at 31 frames per second (FPS) for 550x802 and 18 FPS at 1100x1604 on an NVIDIA RTX3080 graphics card, which includes deformation, rendering, and CNN. When omitting the CNN the speed increases to 43 and 38 FPS, respectively. As a comparison, GHA runs at 22 FPS at 1024x1024 on the same machine. Fig. 8 indicates the number of Gaussians during training, which we limit to a maximum of 250k. We train all our avatars, and baselines, until convergence, which roughly takes 7 hours for GA (on an RTX2080), 30 hours (on an RTX3090) for GHA and our method, and 60 hours (on an RTX2080) for MVP.
Data Preparation
@@ 5.3. Self-Reenactment
Our main evaluation is concerned with the self-reenactment task. For this purpose, all avatars are trained on a set of 21 training sequences alongside their respective tracking results. To evaluate the avatars, they are animated using the tracked expressions from a held-out test sequence. We present qualitative and quantitative results in Fig. 3 and Table 1, respectively, and recommend the reader to consider the supplemental video for temporal results. Our predicted self-reenactments portray the unseen expression more accurately and contain sharper details in relatively static areas like the hair region.
Importance of the Underlying Tracking